OpenStudy (anonymous):

A mass of liquid of density p is mixed with an equal mass of another liquid of density 2p. No change of total volume occurs. What is the density of the liquid mixture? 1) 4p/3 2) 3p/2 3) 5p/3 4) 3p

OpenStudy (paki):

it's option A....

OpenStudy (anonymous):

Can you please explain?

OpenStudy (anonymous):

firstly the mass wont change, so it is the volume of each liquid that changes p= m/v------>1 2p = m / v'-----2 2/1 = v' =0.5v total mass = 2m total volume = v + 0.5v find density =total mass/total volume

OpenStudy (paki):

@emcrazy14... please have a look here, you will get the solution..... https://answers.yahoo.com/question/index?qid=20070610053914AAIPwLo